# Radio Map-Based Spectrum Sharing for Joint Communication and Sensing

### v1

@@ V CONCLUSIONS
In this paper, we have addressed C&S interference within a distributed network. Given that the available C&S cooperation is limited in the real world, we have proposed a radio-map-based C&S cooperation framework. The radio map has been used to estimate the large-scale CSI so that the pilot-based C&S interactions are bypassed. With the large-scale CSI, we have devised a joint power allocation scheme to maximize the radar detection performance while ensuring the ergodic rate requirement. To handle the non-convex C&S metrics, we have employed the techniques of auxiliary-function-based scaling and fractional programming. The proposed problem has been effectively solved using an iterative algorithm. Simulation results have verified that the extrinsic information, i.e., positions and surroundings, is effective to coordinate conflicting signals in the spatial domain. The proposed scheme provides a viable solution for the coexistence of C&S systems in a loose cooperation manner.
@@ Appendix B PROOF OF CONVERGENCE OF THE ITERATIVE ALGORITHM
$f(\mathbf{P}_{c},p_{r_{i}},\beta_{i})=2\beta_{i}\sqrt{[h_{{r_{i}},i}]^{H}p_{r_{i}}h_{{r_{i}},i}}-(\beta_{i})^{2}\sigma^{2}_{r_{i}}(\mathbf{P}_{c}).$

### v2

@@ V CONCLUSIONS
In this paper, we have addressed C&S interference within a distributed network. Given that the available C&S cooperation is limited in the real world, we have proposed a radio-map-based C&S cooperation framework. The radio map has been used to estimate the large-scale CSI so that the pilot-based C&S interactions are bypassed. With the large-scale CSI, we have devised a joint power allocation scheme to maximize the radar detection performance while ensuring the ergodic rate requirement. To handle the non-convex C&S metrics, we have employed the techniques of auxiliary-function-based scaling and fractional programming. The proposed problem has been effectively solved using an iterative algorithm. Simulation results have verified that the extrinsic information, i.e., positions and surroundings, is effective in coordinating conflicting signals. The proposed scheme provides a viable solution for the coexistence of C&S systems in a loosely cooperated manner.
@@ Appendix B PROOF OF CONVERGENCE OF THE ITERATIVE ALGORITHM
$f(\mathbf{P}_{c},p_{r_{i}},\beta_{i})=2\beta_{i}\sqrt{[h_{{r_{i}}\rightarrow{r_{i}}}]^{H}p_{r_{i}}h_{{r_{i}}\rightarrow{r_{i}}}}-(\beta_{i})^{2}\sigma^{2}_{r_{i}}(\mathbf{P}_{c}).$